OpenStudy (anonymous):

The position of a softball tossed vertically upward is described by the equation y=c1(t)−c2(t^2), wherey is in meters,tin seconds,c1=6.71 m/s, and c2= 2.39 m/s^2. How would you find the acceleration at a specific t value?

zepdrix (zepdrix):

\[\Large y=c_1t-c_2t^2\] `Position` is usually denoted by s as a function of t:\[\Large s(t)=c_1t-c_2t^2\] The instantaneous rate of change of postion gives us velocity,\[\Large s'(t)=v(t)\]The second derivative of position, or the first derivative of velocity, gives us acceleration.\[\Large s''(t)=a(t)\] Then to find the acceleration at a specific \(\Large t\) value, let's say at \(\Large b\), we would plug that value into the acceleration function. \[\Large s''(b)=a(b)\]

zepdrix (zepdrix):

So it looks like you'll want to find the second derivative of your function :o Understand how to do that?

OpenStudy (anonymous):

I did that and got 4.78 but apparently that is wrong

zepdrix (zepdrix):

\[\Large s(t)=c_1t-c_2t^2\]\[\Large s'(t)=c_1-2c_2t\]\[\Large s''(t)=-2c_2\] Hmm it looks like you've got the right idea :) Did you remember the negative sign?

OpenStudy (anonymous):

nope! lol

OpenStudy (anonymous):

Always the little things :/

zepdrix (zepdrix):

Hehe yah ^^

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

watcha doubt its tough rite

OpenStudy (anonymous):

do u mind givin me a medal plzz

OpenStudy (anonymous):

watcha doubt its tough rite

OpenStudy (anonymous):

hi